OpenStudy (anonymous):

Can you show me how to solve this equation using substitution 5y+4x=8 y=5x+10

OpenStudy (amistre64):

yes. since it is stated that y = 5x +10; we can just substitute that value for "y" in the other equation and slove for "x"

OpenStudy (amistre64):

5(5x + 10) + 4x = 8 25x + 50 + 4x = 8 29x + 50 = 8 29x = 8 - 50 29x = -42 x = -42/29

OpenStudy (amistre64):

reduce if you can...

OpenStudy (anonymous):

could you show me what I did wrong ill show you how I solved it

OpenStudy (amistre64):

go right ahead :)

OpenStudy (anonymous):

5y+4x=8 y=5x+10 5(5x+10)+4x=8 25x+50+4x=8 21x+50=8 21x=-42 x=-2 y=5(-20)+10 y=-10+10 y=0

OpenStudy (anonymous):

That is what I did

OpenStudy (anonymous):

And I was pretty sure that was the answer.

OpenStudy (amistre64):

25x + 4x does NOT equal 21x

OpenStudy (anonymous):

Oh sorry I made a typing misatke to y=5(-2)+10 should be the one line

OpenStudy (anonymous):

Ithought you subtracted

OpenStudy (amistre64):

nope; when it says 25 + 4, it actually means to add. That little (+) sign means add....

OpenStudy (anonymous):

I know but I was taught that you did the opposite to get rid of it

OpenStudy (amistre64):

why would you want to get rid of it? it is a vital part of the equation. Now, if it was on the OTHER side of the (=) sign, you would subtract it from BOTH sides to move it. But, if it is on the SAME side, it is simply addition.... Does that make sense?

OpenStudy (anonymous):

Look at your original question that you asked. Are you sure it doesn't say -4x in the problem you were given? 5y - 4x = 8?

OpenStudy (anonymous):

Hold on let me check quick

OpenStudy (amistre64):

If you want to move it to the othe side of the (=) sign we do this: 25 + 4 = 29 25 + 4 (-4) = 29 (-4) 25 = 29 - 4 But when it is on the same side, just add it together: 25 + 4 = 29 29 = 29

OpenStudy (anonymous):

5y + 4x = 8 y + 5x = 10 it says ''solving a system of equations by substitution''

OpenStudy (amistre64):

That is a different set of equations that what you give to begin with; please verify which one is correct please...

OpenStudy (anonymous):

Im sorry

OpenStudy (anonymous):

The last one is the correct one

OpenStudy (amistre64):

ok :) What does "y" equal then?: y = 10 - 5x substitute that "value" into the top equation to get: 5(10 - 5x) + 4x = 8 50 - 25x + 4x = 8 -25x + 4x = 8 -50 -21x = -42 x = -42/-21 = 2 Use this solid value of x=2 in the bottom equation to solve for "y" y + 5(2) = 10 y + 10 = 10 y = 10 - 10 y = 0 x = 2

OpenStudy (anonymous):

OK thanks im gonna look over this for a minute

OpenStudy (anonymous):

Im sorry im kinda confussed I was taught in a different way and im trying to figure this all out

OpenStudy (anonymous):

I got y=5x+10

OpenStudy (amistre64):

you gave, for the bottom equation: y + 5x = 10 ; to get "y" all by its little lonesome, we need to subtract (5x) from both sides to keep the same values... y +5x -5x = 10 - 5x y + 0 = 10 - 5x y = 10 - 5x Do you see it?

OpenStudy (anonymous):

Not really im sorry

OpenStudy (amistre64):

tell me where I am losing you then, let me know what is confusing you here. And then I can show you the right steps...

OpenStudy (amistre64):

Can you tell me what the (=) sign means?

OpenStudy (anonymous):

See for Y im somehow got -5x+10

OpenStudy (amistre64):

-5x + 10 is correct, and it is the same thing as 10-5x am i right or wrong?

OpenStudy (anonymous):

Oh so then I did have it right?

OpenStudy (amistre64):

you wrote, and I quote " I got y=5x+10" earlier. If what you meant was "I got y= -5x+10" then yes, you are right...

OpenStudy (anonymous):

Yeah thats what I meant im sorry I cant type today at all

OpenStudy (amistre64):

:) paper and pencil is always easier when doing math

OpenStudy (anonymous):

Yeah it is but thanks for showing me that now im gonna rework the problem and see if i get it right

OpenStudy (anonymous):

so should the set up be 5(-5x+10)+4x=8

OpenStudy (amistre64):

Yes

OpenStudy (anonymous):

Ok I think im getting this down now

OpenStudy (amistre64):

you are doing great :)

OpenStudy (anonymous):

Thanks I have really been having a tough time this year

OpenStudy (anonymous):

Does y = 2

OpenStudy (anonymous):

and x=0

OpenStudy (amistre64):

yes; y=0 and x=2. To double check; plug those values into the original equations to see if they are true.

OpenStudy (anonymous):

Ok thanks a lot :)